If ‘+’ means ‘x’, ‘-‘ means ‘÷’ , ‘x’ means ‘-‘ and ‘÷’ means ‘+’ then what will be the value of
14 ÷ 42 – 6 x 3 + 2
1
Expert's answer
2014-06-04T04:17:24-0400
The expression 14 ÷ 42 – 6 x 3 + 2 would be rewrite as 14 + 42 ÷ 6 - 3 x 2=14+7-6=15.

Answer. 15